Pls Help *MTH282*
Suppose the position vector of X and Y are (1,2,4) and (2,3,5), find the position vector of a point Z that bisect XY in the ratio 2:3
A 7i+12j+22k
B. 7i-12j+22k
C.frac{1}{7} (7i+12j+22k)
D.frac{1}{17} (7i-12j+22k)
1
Expert's answer
2020-05-21T12:22:38-0400

Division of a line segment in a given ratio is: "x=\frac{x_1+ky_1}{1+k},y=\frac{x_2+ky_2}{1+k},z=\frac{x_3+ky_3}{1+k}" where

"(x_1,x_2,x_3)" is"(1,2,4)" , "(y_1,y_2,y_3)" is "(2,3,5)" ,"k=\frac{2}{3}" ,hence "x=7\/5,=12\/5,z=22\/5"

hence Z= "\frac{1}{5}(7i+12j+22k)"